OpenStudy (anonymous):

Determine the solution of the system. Show your work. Y=5x-9 y=2x+6

OpenStudy (danjs):

Set y = y 5x - 9 = 2x + 6

OpenStudy (anonymous):

Huh

OpenStudy (anonymous):

I'm not good at this stuff at all

OpenStudy (danjs):

If y = something, and y also = something else, you can set the something= to something else.

OpenStudy (danjs):

solve 5x - 9 = 2x + 6 for x

OpenStudy (anonymous):

X=3 I'm guessing ?

OpenStudy (anonymous):

-3x

OpenStudy (danjs):

subtract 2x from both sides of the equation first 5x - 2x - 9 = 2x - 2x + 6 3x - 9 = 6 Now add 9 to both sides of the equation, 3x - 9 + 9 = 6 + 9 3x = 15 Divide Both sides of the equation by 3. 3/3 x = 15/3 x = 5

OpenStudy (danjs):

Use that value , x = 5, and put it into either of the original equations for y= , and calculate y when x=5

OpenStudy (danjs):

The solution is the point where the two lines, y = 5x - 9 and y = 2x + 6 intersect on the coordinate plane. (x,y)

OpenStudy (anonymous):

I still don't get it .... 15,1 or 2,1

OpenStudy (anonymous):

@DanJs

OpenStudy (anonymous):

@DanJS

jimthompson5910 (jim_thompson5910):

do you see how Dan got x = 5 ?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

so the answer will look like this: (5, ?) where you replace the question mark with the y coordinate

jimthompson5910 (jim_thompson5910):

you can pick on any equation with x and y in it like y = 5x - 9 and plug in x = 5 to get the value of y

jimthompson5910 (jim_thompson5910):

y = 5x - 9 y = 5(5) - 9 ... replace x with 5 y = ???

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

yes y = 16

jimthompson5910 (jim_thompson5910):

we can use the other equation as well y = 2x + 6 y = 2(5) + 6 ... replace x with 5 y = 10 + 6 y = 16 and we get the same answer for y

jimthompson5910 (jim_thompson5910):

and as Dan is showing, the two lines intersect at (5,16) so that visually confirms things